OpenStudy (anonymous):

What is -8 + 4 is equivalent to?

OpenStudy (texaschic101):

-8 + 4 = - 4

OpenStudy (hunus):

 If you were to move right 4 places from -8 on a number line what number would you arive at

OpenStudy (anonymous):

thank you could u guys help me with a couple more?

OpenStudy (texaschic101):

sure

OpenStudy (anonymous):

Which of the following expressions is equivalent to 3? Answers:7 + (-4) -7 + 4 7 - (-4) -7 - (-4)

OpenStudy (texaschic101):

7 + (-4) = 3

OpenStudy (anonymous):

wow you are smart thanks

OpenStudy (texaschic101):

when the signs are both positive, you add them (even if their both negative. When the signs are different, you subtract and take the sign of the biggest number.

OpenStudy (anonymous):

whoa that sounds familiar thanks

OpenStudy (anonymous):

u are in deed a lifesaver

OpenStudy (texaschic101):

can you do this one.... -4 + (-4) ? remember, the signs are both negative

OpenStudy (anonymous):

oh thats right

OpenStudy (texaschic101):

do you know the answer ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

tell me

OpenStudy (anonymous):

ok -4

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

whoa kinda hard....

OpenStudy (texaschic101):

wrong.... -4 + (-4) = - 8 if the signs are the same, you add them and use that sign. if they are different... -10 + 5 = - 5 (you subtract, and take the sign of the biggest number

OpenStudy (anonymous):

why do i suck so much in math? im great in every other class!!!!!!!!!!!!!!

OpenStudy (anonymous):

ok makes sence...

OpenStudy (texaschic101):

don't fret....math is not my best subject either. You really just need to learn the simple rules... Don't get confused...because multiplying/dividing and adding/subtracting rules are not the same

OpenStudy (texaschic101):

- plus - = - + plus - = - + plus + = +

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

um hold up still computing

OpenStudy (texaschic101):

just remember, if the signs are the same, you add them and use that sign. If the signs are different, you subtract and take the sign of the biggest number.

OpenStudy (anonymous):

ok got it and no more questions

OpenStudy (texaschic101):

you are smart....you can understand this. I wish you luck :)